OpenStudy (muhtasim):

The percentage of research articles in a prominent journal written by researchers in the United States can be modeled by A(t) = 25 + (36/1 + 0.6(0.7)^−t'), where t is time in years (t = 0 represents 1983). Numerically estimate lim as t→+infinity A(t).

OpenStudy (muhtasim):

\[A(t) = 25 + \frac{ 36 }{ 1+0.6(0.7)^{-t} }\]

OpenStudy (muhtasim):

Estimate \[estimate \lim_{x \rightarrow +\infty}A(t) numerically. \]

OpenStudy (irishboy123):

perhaps numerically estimate literally means getting your scientific calculator out and putting some huge numbers in for \(t\), in which forget case about the rest of this .....but note that: \(\lim\limits_{t \to \infty} 0.6(0.7^{-t}) = 0.6 \lim\limits_{t \to \infty} \left( \dfrac{10}{7} \right)^{t} \) And \(\lim\limits_{t \to \infty} 25 + \dfrac{ 36 }{ 1+0.6(0.7)^{-t} } \) \(= 25 + \dfrac{ 36 }{ 1+ \lim\limits_{t \to \infty} 0.6(0.7)^{-t} }\)